OpenStudy (kaylala):

proving identities: (1-3cos-4cos^2)/(sin^2) = (1-4cos)/(1-cos)

OpenStudy (superdavesuper):

u can start by grouping (1-3cos-4cos^2) n use sin^2 = 1 - cos^2 for the bottom part...

OpenStudy (kaylala):

@Loser66 @thomaster @Mertsj @Preetha

OpenStudy (loser66):

I am with superdavesuper.

OpenStudy (kaylala):

i dont get him

OpenStudy (superdavesuper):

try grouping 1-3x-4x^2 into (1....)*(1....)....in this case x represents sin

OpenStudy (mertsj):

\[\frac{1-4\cos x)(1+\cos x)}{(1-\cos x)(1+\cos x)}\]

OpenStudy (mertsj):

Perhaps you can finish it.

OpenStudy (superdavesuper):

thanks @Mertsj unrelated....wooohooo i didnt know LaTex is working again ;)

OpenStudy (mertsj):

Wonderful, isn't it?

OpenStudy (kaylala):

thanks

OpenStudy (superdavesuper):

welcome :)